Simplify: 5(x - 6) + 2x - 10
Found 2 solutions by HyperBrain, madhan_math:
Answer by HyperBrain(694)

5(x - 6) + 2x - 10=(5x-5*6)+2x-10
5(x - 6) + 2x - 10=(5x-30)+2x-10
5(x - 6) + 2x - 10=(5x+2x)+(-30-10)
5(x - 6) + 2x - 10=7x-40
Power up,
HyperBrain!

Answer by madhan_math(34)

5(x-6)+2x-10=???? is your question.
Solution:
Let 5x-30 + 2x-10=0
adding x terms and constants, implies that
7x-40=0
7x=40
x=40/7
or x= 5.71